# FSGe: A fast and strongly-coupled 3D fluid-solid-growth interaction method

### v1

@@ 2.2.2 Equilibrated Constrained Mixture Model
The CMMe directly predicts the long-term evolved state in mechanical and biological equilibrium, i.e., at $s\to\infty$, where production again balances removal. Note that CMMe can still be used to predict transient G&R phenomena as long as the process is considered to be quasi-equilibrated, similar to classical problems in mechanics of quasi-equilibrated problems. Toward this end, it is assumed that the imposed perturbations to the homeostatic state are slow relative to the adaptive G&R process. For this purpose, the G&R process is broken up into several quasi-static load steps where perturbations are applied in increments. This rate-independent approach does not require computing hereditary integrals (7), thus drastically lowering the computational demand. In CMMe, the stimulus function in Equation (5) reduces to
@@
where $\hat{\bm{\sigma}}^{\alpha}_{h}$ are the equilibrated constituent-specific Cauchy stresses, $\bm{\sigma}^{x}$ is the mixture stress, and $p_{h}$ is the equilibrated Lagrange multiplier that enforces mechanobiological equilibrium. We define the IMS invariant throughout this work as
@@
The invariant $\sigma^{x}_{Ih}$ can be explicitly calculated from the mixture stress $\bm{\sigma}^{x}$. The original homeostatic IMS $\sigma_{Io}$ is stored locally from the pre-loading stage. Note that the tangential stiffness matrix of the CMMe material model only has minor symmetries, no major symmetries. This can pose challenges for iterative linear solvers and their preconditioners. We thus use the distributed memory sparse direct linear solver MUMPS [41] within the PETSc suite [42].
@@ 3 Fluid-solid-growth coupling
In classic FSI, the solid material behavior is purely determined by an elastic response to boundary conditions and interface conditions at the fluid-solid interface $\Gamma_{\text{int}}$. In contrast, the CMMe material behavior is characterized by a long-term mechanobiological response to mechanical stimuli. In this section, we first highlight key similarities and differences between FSI and FSGe problems and then derive a robust coupling scheme. WSS is a crucial stimulus in the models used in this work. Its local calculation is a key difference between purely solid G&R models and FSGe:
@@
In the G&R model, local WSS is approximated with the Poiseuille flow-derived WSS (1). The evolved local inner radius $a_{h}$ of the blood vessel is approximated locally at each Gauß point of the finite elements. The approximation uses the radial coordinate $r_{o}$ and inner radius $a_{o}$ in original homeostasis as well as the evolved local circumferential stretch $\lambda_{\theta}$ and radial stretch $\lambda_{r}$ with unit vectors $\mathbf{e}_{i}$ [5]. In the FSGe model, we use the magnitude of the local WSS vector $\bm{\tau}_{w}$ extracted at the fluid-solid interface from a steady-state fluid solution. In response to local changes in WSS, the endothelium alters its production of vasoactive molecules that rapidly diffuse into the wall and cause vascular smooth muscle cell relaxation or contraction [43]. Instead of solving the diffusion equation, this work assumes that the WSS stimulus at the interface $\Gamma_{\text{int}}$ effects each radial location equally within the vessel wall. Thus, the FSGe model is coupled via
@@
to the solid stress tensor $\bm{\sigma}^{\text{s}}$. Modeling the effect of vasoactive molecules, at everly location in the vessel the solid stress $\bm{\sigma}^{\text{s}}$ depends on WSS $\bm{\tau}_{w}$, which, in turn, depends on the fluid stress tensor $\bm{\sigma}^{\text{f}}_{\Gamma_{\text{int}}}$ at the interface $\Gamma_{\text{int}}$. This interaction differs from classic FSI problems, where fluid and solid are only coupled at the interface. In FSGe, the fluid-solid interface is coupled to the whole solid volume. Furthermore, in blood flow, WSS $\tau_{w}$ is typically five orders of magnitude smaller than the IMS $\sigma_{I}$. Thus, WSS has a negligible mechanical influence in vascular FSI models, but, depending on the gain ratio $K_{\tau\sigma}$, a large biological influence in the FSGe model. Also, solid G&R takes place on a time scale $T^{\text{s}}\gg T^{\text{f}}$ of weeks or months, which is much larger than the fluid dynamics time scale $T^{\text{f}}$. Thus, we can derive the interface velocity from the interface displacements $\bm{d}$ as
@@ 3.1 Monolithic coupling
where $\partial_{\mathbf{U}}\,\mathbf{f}$ denotes the partial derivative of $\mathbf{f}$ with respect to $\mathbf{U}$. All blocks in Equation (22) are treated in the same mathematical framework and implemented in svFSIplus. For FSI problems, the advantages of monolithic coupling can be faster convergence and higher accuracy than the partitioned approach. However, it may require more computational resources and a more intrusive software implementation. Independent of these challenges, we found a monolithic coupling approach to be intractable for an FSGe problem. Solving Equation (22) requires solving both fluid and solid fields with the same time step size, typically $T_{\text{FSI}}\sim 1\,\text{ms}$ or below in cardiovascular FSI problems [47]. However, as discussed in Section 3, solid G&R takes place over a much longer time scale than blood flow fluid dynamics. In fact, the CMMe returns the mechanobiological equilibrated G&R response after an infinitely long period $s\to\infty$ at every load step $t$. This limit renders solving Equation 22 with a uniform fluid-solid time step size intractable, and we, therefore, chose a partitioned scheme for FSGe. Note that we refer to $t$ throughout this work as a quasi-static G&R load step opposed to a time step in a dynamic problem. All load steps represent a long-term equilibrium at $s\to\infty$, typically for different insult severities.
@@ 3.2 Partitioned coupling
with a stiffness matrix $\mathbf{K}$. Inspired by Degroote [45], we outline the steps within the fluid and solid solvers in Algorithms 1 and 2, respectively. Partitioned FSI offers the advantage of reusing and combining existing fluid and solid solvers with solution techniques targeted to the respective domain. Note that while independent solvers could be used for $\mathcal{F}$ and $\mathcal{S}$, we use our open-source multi-physics solver svFSIplus for both [25]. Crucially, this allows us to natively split both time scales arising in the FSGe problem. We thus focus on a partitioned coupling approach for the remainder of this work.
@@
2:Propagate WSS through solid domain
@@ 3.2.2 Strong coupling
is smaller than a given tolerance $\epsilon_{0}$. The most basic form of strong coupling is the Gauß-Seidel iteration scheme depicted in line 8 of Algorithm 4. Here, fluid mesh deformation, fluid, and solid are executed sequentially, and we use the output from the previous coupling iteration as the input for the current iteration:
@@
The strongly coupled FSGe method is outlined in Algorithm 4. We use a Gauß-Seidel scheme exclusively at $t=0$ for the pre-loading stage in CMMe [5] in line 8. Here, the displacements are negligibly small, and there is only weak coupling as pre-loading does not depend on local WSS. The G&R stage in CMMe at $t>0$ requires the more robust IQN-ILS coupling approach in line 12. Note that the performance of the algorithm also crucially depends on an accurate prediction of displacements $\mathbf{d}_{t,0}$ at the beginning of each load step. If a previous load step is available, we use the linear predictor in Algorithm 5. Further, note that at the beginning of each coupling iteration $k=1$, we obtain input displacements $\mathbf{d}_{t,k}$ from a static relaxation step (30). Similarly, we use static relaxation for the first five coupling iterations at $t=1$, i.e., the first G&R load step, to collect an initial number of iterations for IQN-ILS.
@@ 4 Illustrative computational examples
Diagonal collagen orientation $\alpha_{0}$ 29.9°
@@ 4.3 Fluid dynamics
Figure 12 shows velocity (left) and pressure (right) for the vessel’s longitudinal midsection for different gain ratios $K_{\tau\sigma,o}$ in the FSGe. Note that, due to the absence of local fluid dynamics in the G&R model, we cannot compare the models. The geometries correspond to the right column in Figure 5. Recalling the fluid boundary conditions from Section 4, it is evident that all flows have identical (parabolic) inflow conditions on the very left and identical outflow pressures of 104.9 mmHg on the very right. While there is little variation in pressure fields for the sampled gain ratios $K_{\tau\sigma,o}$, there are some notable differences in the velocity field. There is a small recirculation zone visible at the top apex of the aneurysm. Due to the incompressibility of blood, for $K_{\tau\sigma,o}=0.0$ (top), the doubling in radius inside the aneurysm drastically decreases local velocity. This effect is less pronounced as the maximum radius decreases with increasing gain ratio $K_{\tau\sigma,o}$. Interestingly, as the gain ratio approaches $K_{\tau\sigma,o}\to 1$, the velocity is deflected upward into the top part upstream of the aneurysm.
@@ 5 Discussion
Latorre and Humphrey [5] introduced the CMMe as a fast, rate-independent version of the constrained mixture model (CMM). It directly yields the long-term equilibrated G&R response, i.e., at $s\to\infty$. The computational cost of CMMe is comparable to a solid mechanics FEM solution with a hyperelastic material. Together with our implicit coupling scheme, this makes the FSGe highly efficient. On the other hand, it may be critical to model transient G&R in cases where the insult is applied slowly compared to the G&R response. For example, tissue-engineered vascular grafts (TEVGs) in Fontan patients can develop a severe stenosis, which sometimes spontaneously resolves [14, 15]. Here, the time-resolved development of G&R is crucial to studying the mechanisms behind the stenosis formation and resolution [13], which can be modeled with FSGe [12]. However, we can only study phenomena that will yield a long-term mechanobiological equilibrium. Yet, many clinically interesting applications involve G&R processes that do not necessarily yield an evolved equilibrial state [57]. For example, theories have been developed to quantify the stability or instability of G&R processes such as aneurysm formation [58, 59]. Thus, transient G&R models like the one developed by Schwarz et al. [13] will be necessary to study fluid-solid-growth interaction (FSG) in unstable cases.
@@ 5.4 Definition of original homeostasis
In computational biomechanics, the solid reference configuration, often obtained from in vivo medical imaging, is typically not stress-free due to fluid or other loads. This requires either finding a stress-free reference configuration, which might not exist [64], or local residual stresses so that the initial configuration is in equilibrium with fluid loads. Many strategies have been developed to address this issue [65, 66, 67]. In G&R, we need to satisfy a biological equilibrium, i.e., original homeostasis $o$, in addition to the aforementioned mechanical equilibrium. In this work, we achieve a mechanobiological equilibrated original homeostasis by solving a non-linear equation for a thin-walled cylinder for the initial in vivo blood pressure $P_{o}$. Here, the inner and outer radii are fixed, as well as the homeostatic parameters in Table 2. This approach works well in a given ideal cylinder with an unknown uniform in vivo blood pressure that can be freely chosen in the model. However, it cannot be generalized to in vivo blood vessel geometries that were imaged under measured blood pressure. Given the limited prior work with 3D CMM, there are few references for pre-loading in vivo G&R models. The algorithm proposed by Wu et al. [68] solves a minimization problem by locally varying fiber alignment. Laubrie et al. [69] prescribed a transmurally varying prestretch gradient for constituents to establish an initial homeostatic state. Gebauer et al. [70] introduced an iterative algorithm to prestress a fixed patient-specific cardiac geometry under fixed ventricular pressures. The algorithm solves for the local isotropic elastin prestretch until the model is in mechanobiological equilibrium up to a specified displacement tolerance. The choice of elastin prestretch was motivated by the fact that elastin is the only constituent that is assumed to not turn over in the G&R time scale considered, given its long half-life [71]. In the future, this approach could be used to define the original homeostasis in FSGe problems by either locally solving for the elastin prestretch or other quantities, such as vessel thickness.
@@ 5.5 Influence of local wall shear stress
$\tau_{w}\sim Q\cdot\frac{1}{r^{3}},\quad\sigma_{\theta}\sim P\cdot\frac{r}{h},\quad\sigma_{r}\sim-P,\quad\sigma_{z}\sim F\cdot\frac{1}{rh},$ (38)
with radius $r$, wall thickness $h$, flow rate $Q$, pressure $P$, and axial force $F$. We specifically developed our FSGe coupling method to not rely on Poiseuille flow and thin-walled assumptions but rather the numerical solution of the 3D Navier-Stokes equations and 3D finite strain elasticity, respectively. Yet, we studied aneurysmal development under steady-state flow with a low Reynolds number of $\text{Re}\approx 170$ which resulted in a flow profile similar to parabolic even in aneurysmal regions of the blood vessel. Thus, comparing the fluid and solid field solutions derived from our FSGe model to these simplifying assumptions gives us insights into the feedback loop of biomechanically regulated variables (Figure 1) and helps us explain the differences in mechanobiological equilibria. In our examples, the flow rate $Q$ was constant, and there was little change in local pressure $P$ throughout the fluid domain. We can thus derive the following relationships for the change in WSS and intramural stimuli
$\Delta\tau_{w}$ $\sim\frac{r_{o}^{3}}{r_{h}^{3}}-1,$ (39)
$\Delta\sigma_{I}$ $\sim\frac{h_{o}}{r_{o}}\frac{r_{h}}{h_{h}}-1.$ (40)
The differences in aneurysmal growth were associated with these different definitions of stimuli and, therefore, different values of mechanobiological equilibria in the two models. Our illustrative computational examples highlight the importance of incorporating local fluid-derived WSS in the FSGe model. In the FSGe model, the fluid couples the deformation in the aneurysmal top part of the blood vessel with the healthy bottom part. In the purely solid G&R model, the WSS in these regions is evaluated independently from local deformations. The stimuli $\Delta\tau_{w}<0$ and $\Delta\sigma_{I}<0$ were downregulated to a much larger extent in FSGe than in G&R at the bottom of the aneurysm. In the G&R model, relying on relationship (39) resulted in an increase in evolved radius $r_{h}$ and thickness $h_{h}$. In the FSGe model, the fluid-derived WSS, $\Delta\tau_{w}$ decreased with decreasing $r_{h}$ which is contrary to (39). This was possible due to the Navier-Stokes-derived WSS (Figure 11, right) and the upward deflection of the blood flow inside the aneurysm (Figure 12). In both models, this change was accompanied by an increase in thickness $h_{h}$ according to (40).

### v2

@@ 2.2.2 Equilibrated Constrained Mixture Model
The CMMe directly predicts the long-term evolved state in mechanical and biological equilibrium, i.e., at $s\to\infty$, where production again balances removal. Note that CMMe can still be used to predict transient G&R phenomena as long as the process is considered to be quasi-equilibrated, similar to classical problems in mechanics of quasi-equilibrated problems. Toward this end, it is assumed that the imposed perturbations to the homeostatic state are slow relative to the adaptive G&R process. For this purpose, the G&R process is broken up into several quasi-static load steps $t$ where perturbations are applied in increments, as opposed to a time step in a dynamic problem. All load steps represent a long-term equilibrium at $s\to\infty$, typically for different insult severities. This rate-independent approach does not require computing hereditary integrals (7), thus drastically lowering the computational demand. In CMMe, the stimulus function in Equation (5) reduces to
@@
where $\hat{\bm{\sigma}}^{\alpha}_{h}$ are the equilibrated constituent-specific Cauchy stresses, $\bm{\sigma}^{x}$ is the extra part of the mixture stress, and $p_{h}$ is the equilibrated Lagrange multiplier that enforces mechanobiological equilibrium. We define the IMS invariant throughout this work as
@@
The invariant $\sigma^{x}_{Ih}$ can be explicitly calculated from the mixture stress $\bm{\sigma}^{x}$. The original homeostatic IMS $\sigma_{Io}$ is stored locally from the pre-loading stage. Note that the tangent stiffness matrix of the CMMe material model only has minor symmetries, no major symmetries. This can pose challenges for iterative linear solvers and their preconditioners. We thus use the distributed memory sparse direct linear solver MUMPS [49] within the PETSc suite [50].
@@ 3 Fluid-solid-growth coupling
In classic fluid-structure interaction (FSI), the solid material behavior is determined by an elastic response to boundary conditions and interface conditions at the fluid-solid interface $\Gamma_{\text{int}}$. In contrast, the CMMe material behavior is characterized by a long-term mechanobiological response to mechanical stimuli. In this section, we first highlight key similarities and differences between FSI and FSGe problems and then derive a robust coupling scheme. WSS is a crucial stimulus in the models used in this work. Its local calculation is a key difference between purely solid G&R models and FSGe:
@@
In the G&R model, local WSS is approximated with the Poiseuille flow-derived WSS (1). The evolved local inner radius $a_{h}$ of the blood vessel is approximated locally at each Gauß point of the finite elements. The approximation uses the radial coordinate $r_{o}$ and inner radius $a_{o}$ in original homeostasis as well as the evolved local circumferential stretch $\lambda_{\theta}$ and radial stretch $\lambda_{r}$ with unit vectors $\mathbf{e}_{i}$ [5]. In the FSGe model, we use the magnitude of the local WSS vector $\bm{\tau}_{w}$ extracted at the fluid-solid interface from a steady-state fluid solution. In response to local changes in WSS, the endothelium alters its production of vasoactive molecules that rapidly diffuse into the wall and cause vascular smooth muscle cell relaxation or contraction [51]. Instead of solving the diffusion equation, this work assumes that the WSS stimulus at the interface $\Gamma_{\text{int}}$ affects each radial location equally within the vessel wall. Thus, the FSGe model is coupled via
@@
to the solid stress tensor $\bm{\sigma}^{\text{s}}$. By modeling the effect of vasoactive molecules at every location in the vessel, the solid stress $\bm{\sigma}^{\text{s}}$ depends on WSS $\bm{\tau}_{w}$, which, in turn, depends on the fluid stress tensor $\bm{\sigma}^{\text{f}}_{\Gamma_{\text{int}}}$ at the interface $\Gamma_{\text{int}}$. This interaction differs from classic FSI problems, where fluid and solid are only coupled at the interface. In FSGe, the fluid-solid interface is coupled to the whole solid volume. Furthermore, in blood flow, WSS $\tau_{w}$ is typically five orders of magnitude smaller than the IMS $\sigma_{I}$. Thus, WSS has a negligible mechanical influence in vascular FSI models, but, depending on the gain ratio $K_{\tau\sigma}$, a large biological influence in the FSGe model. Also, solid G&R takes place on a time scale $T^{\text{s}}\gg T^{\text{f}}$ of weeks to months, which is much larger than the fluid dynamics time scale $T^{\text{f}}$ of milliseconds to seconds. Thus, we can derive the interface velocity from the interface displacements $\bm{d}$ as
@@ 3.1 Monolithic coupling
where $\partial_{\mathbf{U}}\,\mathbf{f}$ denotes the partial derivative of $\mathbf{f}$ with respect to $\mathbf{U}$. All blocks in Equation (22) are treated in the same mathematical framework and implemented in svFSIplus. For FSI problems, the advantages of monolithic coupling can be faster convergence and higher accuracy than the partitioned approach. However, it may require more computational resources and a more intrusive software implementation. Independent of these challenges, we found a monolithic coupling approach to be intractable for an FSGe problem. Solving Equation (22) requires solving both fluid and solid fields with the same time step size, typically $T_{\text{FSI}}\sim 1\,\text{ms}$ or below in cardiovascular FSI problems [55]. However, as discussed in Section 3, solid G&R takes place over a much longer time scale than blood flow dynamics. In fact, the CMMe returns the mechanobiological equilibrated G&R response after an infinitely long G&R period $s\to\infty$ at every load step $t$. This limit renders solving Equation 22 with a uniform fluid-solid time step size intractable, and we, therefore, chose a partitioned scheme for FSGe.
@@ 3.2 Partitioned coupling
with a stiffness matrix $\mathbf{K}$. Inspired by Degroote [53], we outline the steps within the fluid and solid solvers in Algorithms 1 and 2, respectively. Partitioned FSI offers the advantage of reusing and combining existing fluid and solid solvers with solution techniques targeted to the respective domain. Note that while independent solvers could be used for $\mathcal{F}$ and $\mathcal{S}$, we use our open-source multi-physics solver svFSIplus for both [33]. Crucially, this allows us to naturally split both time scales arising in the FSGe problem. We thus focus on a partitioned coupling approach for the remainder of this work.
@@
2:Propagate WSS stimulus through solid domain
@@ 3.2.2 Strong coupling
is smaller than a given tolerance $\epsilon_{0}$. Note that for a sufficiently small $\epsilon_{0}$, the converged solution of the strongly coupled partitioned scheme is identical to the solution of a monolithic scheme. The most basic form of strong coupling is the Gauß-Seidel iteration scheme depicted in line 8 of Algorithm 4. Here, fluid mesh deformation, fluid, and solid are executed sequentially, and we use the output from the previous coupling iteration as the input for the current iteration:
@@
The strongly coupled FSGe method is outlined in Algorithm 4. We use a Gauß-Seidel scheme exclusively at $t=0$ for the pre-loading stage in CMMe [5] in line 8. Here, the displacements are negligibly small, and there is only weak coupling as pre-loading does not depend on local WSS. The G&R stage in CMMe for all $t>0$ requires the more robust IQN-ILS coupling approach in line 12. Note that the performance of the algorithm also crucially depends on an accurate prediction of displacements $\mathbf{d}_{t,0}$ at the beginning of each load step. If a previous load step is available, we use the linear predictor in Algorithm 5. Further, note that at the beginning of each coupling iteration $k=1$, we obtain input displacements $\mathbf{d}_{t,k}$ from a static relaxation step (30). Similarly, we use static relaxation for the first five coupling iterations at $t=1$, i.e., the first G&R load step, to collect an initial number of iterations for IQN-ILS.
@@ 4 Illustrative computational examples
Diagonal collagen orientation $\alpha_{0}$ $29.9^{\circ}$
@@ 4.3 Fluid dynamics
Figure 12 shows velocity (left) and pressure (right) for the vessel’s longitudinal midsection for different gain ratios $K_{\tau\sigma,o}$ in the FSGe. Due to the absence of local fluid dynamics in the G&R model, we cannot compare the models. The geometries correspond to the right column in Figure 5. Recalling the fluid boundary conditions from Section 4, it is evident that all flows have identical (parabolic) inflow conditions on the very left and identical outflow pressures of 104.9 mmHg on the very right. While there is little variation in pressure fields for the sampled gain ratios $K_{\tau\sigma,o}$, there are some notable differences in the velocity field. There is a small recirculation zone visible at the top apex of the aneurysm. Due to the incompressibility of blood, for $K_{\tau\sigma,o}=0.0$ (top), the doubling in radius inside the aneurysm drastically decreases local velocity. This effect is less pronounced as the maximum radius decreases with increasing gain ratio $K_{\tau\sigma,o}$. Interestingly, as the gain ratio approaches $K_{\tau\sigma,o}\to 1$, the velocity is deflected upward into the top part upstream of the aneurysm.
@@ 5 Discussion
Latorre and Humphrey [5] introduced the CMMe as a fast, rate-independent version of the constrained mixture model (CMM). It directly yields the long-term equilibrated G&R response, i.e., at $s\to\infty$. The computational cost of CMMe is comparable to a solid mechanics FEM solution with a hyperelastic material. Together with our implicit coupling scheme, this makes the FSGe highly efficient. On the other hand, it may be critical to model transient G&R in cases where the insult is applied slowly compared to the G&R response. For example, tissue-engineered vascular grafts in Fontan patients can develop a severe stenosis, which sometimes spontaneously resolves [12, 13]. Here, the time-resolved development of G&R is crucial to studying the mechanisms underlying the stenosis formation and resolution [30], which can be modeled with FSGe [28]. However, we can only study phenomena that will yield a long-term mechanobiological equilibrium. Yet, many clinically interesting applications involve G&R processes that do not necessarily yield an evolved equilibrial state [64]. For example, theories have been developed to quantify the stability or instability of G&R processes such as aneurysm formation [65, 66]. Thus, transient G&R models like the one developed by Schwarz et al. [30] will be necessary to study fluid-solid-growth interaction (FSG) in unstable cases.
@@ 5.4 Definition of original homeostasis
In computational biomechanics, the solid reference configuration, often obtained from in vivo medical imaging, is typically not stress-free due to fluid or other loads. This requires either finding a stress-free reference configuration, which might not exist [71], or local residual stresses so that the initial configuration is in equilibrium with fluid loads. Many strategies have been developed to address this issue [72, 73, 74]. In G&R, we also need to satisfy a biological equilibrium, i.e., original homeostasis $o$, in addition to the aforementioned mechanical equilibrium. In this work, we achieve a mechanobiological equilibrated original homeostasis by solving a nonlinear equation for a thin-walled cylinder for the initial in vivo blood pressure $P_{o}$. Here, the inner and outer radii are fixed, as well as the homeostatic parameters in Table 2. This approach works well in a given ideal cylinder with an unknown uniform in vivo blood pressure that can be freely chosen in the model. However, it cannot be generalized to in vivo blood vessel geometries that were imaged under measured blood pressure. Given the limited prior work with 3D CMM, there are few references for pre-loading in vivo G&R models. The algorithm proposed by Wu et al. [75] solves a minimization problem by locally varying fiber alignment. Laubrie et al. [76] prescribed a transmurally varying prestretch gradient for constituents to establish an initial homeostatic state. Gebauer et al. [77] introduced an iterative algorithm to prestress a fixed patient-specific cardiac geometry under fixed ventricular pressures. The algorithm solves for the local isotropic elastin prestretch until the model is in mechanobiological equilibrium up to a specified displacement tolerance. The choice of elastin prestretch was motivated by the fact that elastin is the only constituent that is assumed to not turn over in the G&R time scale considered, given its long half-life [78]. In the future, this approach could be used to define the original homeostasis in FSGe problems by either locally solving for the elastin prestretch or other quantities, such as vessel thickness.
@@ 5.5 Influence of local wall shear stress
$\tau_{w}\sim Q\cdot\frac{1}{{\color[rgb]{0,0,0}a}^{3}},\quad\sigma_{\theta}\sim P\cdot\frac{{\color[rgb]{0,0,0}a}}{h},\quad\sigma_{r}\sim-P,\quad\sigma_{z}\sim F\cdot\frac{1}{{\color[rgb]{0,0,0}a}h},$ (38)
with luminal radius $a$, wall thickness $h$, flow rate $Q$, pressure $P$, and axial force $F$. We specifically developed our FSGe coupling method to not rely on Poiseuille flow and thin-walled assumptions but rather the numerical solution of the 3D Navier-Stokes equations and 3D finite strain elasticity, respectively. Yet, we studied aneurysmal development under steady-state flow with a low Reynolds number of $\text{Re}\approx 170$ which resulted in a flow profile similar to parabolic even in aneurysmal regions of the blood vessel. Thus, comparing the fluid and solid field solutions derived from our FSGe model to these simplifying assumptions gives us insights into the feedback loop of biomechanically regulated variables (Figure 1) and helps us explain the differences in mechanobiological equilibria. In our examples, the flow rate $Q$ was constant, and there was little change in local pressure $P$ throughout the fluid domain. We can thus derive the following relationships for the change in WSS and intramural stimuli
$\Delta\tau_{w}$ $\sim\frac{{\color[rgb]{0,0,0}a}_{o}^{3}}{{\color[rgb]{0,0,0}a}_{h}^{3}}-1,$ (39)
$\Delta\sigma_{I}$ $\sim\frac{h_{o}}{{\color[rgb]{0,0,0}a}_{o}}\frac{{\color[rgb]{0,0,0}a}_{h}}{h_{h}}-1.$ (40)
The differences in aneurysmal growth were associated with these different definitions of stimuli and, therefore, different values of mechanobiological equilibria in the two models. Our illustrative computational examples highlight the importance of incorporating local fluid-derived WSS in the FSGe model. In the FSGe model, the fluid couples the deformation in the aneurysmal top part of the blood vessel with the healthy bottom part. In the purely solid G&R model, the WSS in these regions is evaluated independently from local deformations. The stimuli $\Delta\tau_{w}<0$ and $\Delta\sigma_{I}<0$ were downregulated to a much larger extent in FSGe than in G&R at the bottom of the aneurysm. In the G&R model, relying on relationship (39) resulted in an increase in evolved radius ${\color[rgb]{0,0,0}a}_{h}$ and thickness $h_{h}$. In the FSGe model, the fluid-derived WSS, $\Delta\tau_{w}$ decreased with decreasing ${\color[rgb]{0,0,0}a}_{h}$ which is contrary to (39). This was possible due to the Navier-Stokes-derived WSS (Figure 11, right) and the upward deflection of the blood flow inside the aneurysm (Figure 12). In both models, this change was accompanied by an increase in thickness $h_{h}$ according to (40).